# Kramer’s Escape Rate and Phase Transition Dynamics in AdS Black Holes with Dark Structures

### v2

@@ 3.0.1 Behavioral comparison of landscape free energy
Before any calculations, it is better first compare the various forms of free energy definitions, namely $G_{L}$, U, and L. Theoretically, from the canonical perspective the number of particles and mass should remain constant, consequently only the entropy-temperature and pressure-volume terms will play a role in the first law of thermodynamics. Additionally, given that in Landau’s free energy, $X$ can be considered equivalent to volume, it appears that ultimately, all three introduced forms for the free energy landscape must coincide. In Figs (2a) and (2b), for the model under study, we have plotted these three definitions once according to their general structure and once based on their dimensionless state. Practically, in terms of the shape structure, all three forms have the same trend, but there are also some visible differences, some of which may be attributed to computational precision. Moreover, since Landau’s free energy has dimensions of energy, we used the $G_{c}$ to render it dimensionless, which seems to have increased the degree of displacement, although the shape has been preserved. An intriguing point that comes to mind before concluding this discussion is that in statistical ensembles, such as the grand canonical ensemble where terms related to the chemical potential play a more prominent role, the complexities and structural differences of these three definitions may become more apparent compared to this scenario. Concerning the above discussion and as seen in Fig (2), all the shapes of the free energy landscape (with a little tolerance for displacement) will practically have the same shape structure in the canonical view. Accordingly, the choice and use of each do not seem to have an advantage over the others. For this reason, we will henceforth utilize them under the designation of the free energy landscape, without loss of generality.
Figure 2: Comparison of the Gibbs landscape free energy ($G_{L}$), thermal potential ($U$), and Landau free energy ($L$) in the diagram in two states, i.e., with dimension in Fig (2a) and dimensionless in Fig (2b)
Now, to be able to interpret the behavior of the energy function, we need to have the graph of temperature changes Fig (3a) and Gibbs free energy in terms of temperature Fig (3b).
@@
To better interpret the behavior of free energy, we initially turn to the swallowtail Fig (3b). In this figure, the line (A-N) represents the small black hole, and( C-J ) represents the large black hole, both of which are locally stable with positive heat capacity, and the line (N-C) with negative heat capacity represents the intermediate black hole, which is unstable. As a classical principle, we know that systems always tend to choose states with the least amount of energy. This principle also holds for the free energy landscape, meaning the system will move towards its minimum. In other words, the minima of the energy function correspond to stable states, and its maxima represent unstable states. Considering the overall shape in Fig (4), it is also important to note that in a system with multiple local minima, each of these minima, although locally stable relative to their immediate vicinity, may act as unstable states globally relative to the primary minimum. Even if the system is initially in these states, over time it will progress towards the ultimate global minimum.
In Fig (4a), as is evident, the potential at point A has a global minimum, which belongs to the branch of the small black hole. Over time, with the increase in temperature, in Fig (4b), the local extrema B (minimum) and C (maximum) begin to form. However, they are still in the form of globally unstable states, and the system tends to maintain its position at the primary minimum on the branch of the small black hole. The situation in Fig (4c) is similar to (4b), with the difference that the local extrema have fully formed. With a gradual increase in temperature, the system progresses towards the formation of a large and stable black hole, as shown in Fig (4d). Yet, the small black hole also exists in a stable state, indicating a situation where two equivalent global minima are created. This state corresponds to the isotherm line (H-I) in Fig (3b), which seemingly has a maximum on the branch of the intermediate black hole (I) and a global minimum (H). However, in reality, (H) is located on both the small and large black holes, meaning that a state is created where both the small and large black holes appear simultaneously and stably. In other words, a coexistence phase between the two stable black holes is formed. Further, the process of temperature increase will strengthen the large black hole and weaken the small black hole, Fig(4e) and (4f). The point that can be mentioned at the end of this discussion is related to the second-order phase transition. When we study the system at a temperature higher than the critical temperature, Fig(4g), the system will follow a path similar to (I-H-J) in Fig (3b), and practically it is directly and continuously transferred from the small black hole to the large one.
@@
3.0.2 Kramer’s escape rate
@@
In Fig(5), we approach a state at the onset of phase transition where the chaotic process of phase transition has formed due to primary reasons, meaning the $\gamma$ minimum has found its initial form (as shown in Fig(5b)), although, it has not yet become a global minimum. The small black hole (in the local minimum potential $\alpha$) has begun its intense movement towards the large black hole, considering the formation of a secondary minimum and prevailing conditions, which is clearly observed in Fig(5c). The probability of transition $\alpha\rightarrow\gamma$ is significantly higher than the other transition, Fig(5c), which is why the dominant transfer seems to be entirely from $\alpha\rightarrow\gamma$. Since we are at the beginning of this transition, it seems that the big black hole has not yet formed, or in other words, it lacks enough particles for the formation of reverse processes. Therefore, as it is evident from the Fig(5d), the structure is under the complete control of the transition from a small black hole to a large one, and $\Delta r_{k}$ is completely positive.
@@
Figure 5: t versus x and the First order Phase Transition region (FPT) in Fig (5a),  (5b): The graph of $u(r)$ against x for free parameters and the FPT points,   (5c): comparing $r_{k}$ with respect to $x$ for($\alpha\rightarrow\gamma$) and ($\gamma\rightarrow\alpha$) and the coordinate of contact point   Fig (5d): $\Delta r_{k}$ ($r_{k}(\alpha\rightarrow\gamma)-r_{k}(\gamma\rightarrow\alpha)$) with respect to $x$
In the subsequent stage, we examine a state midway through the phase transition, where the minimum $\gamma$ has assumed a relatively complete form (as depicted in Fig(6a)), and although it has not yet become a global minimum, it is on par in depth with the alpha minimum. If we were to consider this statically and in a single-frame format, it would appear that the escape rate and the probability of movement from the local beta maximum in both directions (i.e., towards the small or the large black hole) should be equal. However, it is important to note that on one hand, the environmental conditions imposing the phase transition on the system, and on the other hand, the shrinking radius —which is somewhat contrary to normal physical laws— still, make the probability of transition $\alpha\rightarrow\gamma$ greater than the other transition, as shown in Fig(6b). For this reason, the dominant transfer seems to continue to be from $\alpha\rightarrow\gamma$. Nevertheless, as is evident in Fig (6b), a collision point has formed between the escape rate diagrams, and beyond this contact point, it is clear that the reverse process, albeit in a very small region at the end of the diagram, predominates over the direct process. It appears that as we approach the end of the phase transition, a negative region (Fig (6c,6d)) is forming which somehow, by reinforcing feedback processes, sets the stage for a change in the initial conditions and ultimately halts the phase transition.
@@
Figure 6: (6a): The $u(r)$ against x for free parameters and the FPT points,   (6b): comparing $r_{k}$ with respect to $x$ for($\alpha\rightarrow\gamma$) and ($\gamma\rightarrow\alpha$) and the coordinate of contact point   (6c):$\Delta r_{k}$ with respect to $x$   (6d): The enlarged detail of the right end of the Fig (6c)
@@
Ultimately, we turn our attention to a state near the end of the phase transition pathway, where the $\gamma$ has taken on a relatively complete shape (as illustrated in Fig(7a)) and has become a global minimum. Although the environmental conditions that are imposing the phase transition on the system continue to make the probability of transition $\alpha\rightarrow\beta$ greater than the other transitions, as depicted in Fig(7b). However, as can be seen from Fig (7b), the coordinates of the contact point have moved significantly upwards, to the extent that the probability of the reverse process cannot be ignored in comparison to the direct process, and now a larger region is dominated by this process, as shown in Fig(7c). As we observed throughout this sequence, the closer we got to the end of the phase transition, the contact point advanced towards, increasing the probability of the reverse transition. It seems that this reverse process may, in some way, be the system’s reaction to prevent the phase transition process from becoming uncontrollable and the radial growth that might overshadow the stability of the black hole. It may also be hypothesized that the black hole, to prevent the phase transition process from becoming uncontrollable, instead of a direct and sudden transition from the small black hole to the large black hole (akin to a completely damped oscillation), performs this process in the form of a slowly damped, quasi-oscillatory manner.
@@
4 4D AdS Einstein-gauss-bonnet-Yang-Mills black hole with a cloud of strings
@@
Since we tried to explain the contents as much as possible in the previous sections, considering the computational similarities, in this section, we will only mention the basic relationships, and we will bring the similar and unnecessary diagrams in Appendix A. The second selected model is the (EGB-YM-CS) black hole. So the metric of 4D AdS Einstein-gauss-bonnet-yang-mills black hole with a cloud of strings is as follows[64],
@@
Figure 7: (7a): The $u(r)$ against x for free parameters and the FPT points,   (7b): comparing $r_{k}$ with respect to $x$ for($\alpha\rightarrow\gamma$) and ($\gamma\rightarrow\alpha$) and the coordinate of contact point   (7c):$\Delta r_{k}$ with respect to $x$
For the main quantities of this model i.e. mass $M$, Hawking temperature $T_{H}$, entropy $S$, pressure $P$ and volume $V$ we will have[64],
@@
where q is the charge of the black hole, ”a” is related to the density of clouds of strings and comes from the energy-momentum tensor of the clouds of strings and $g$ 666 Gauss-Bonnet coupling constant is displayed with g instead of $\alpha$ to avoid mistakes with clouds of strings parameter is a positive Gauss-Bonnet coupling constant.
4.0.1 The landscape free energy
According to equations (2), (3), (4) and (7) for Gibbs and Landau free energy function and thermal potential of this model, we will have,
@@ 4.0.1 The landscape free energy
In this model, we still use relations (17) to make the equations dimensionless.
4.0.2 Kramer’s escape rate
@@
Figure 8: t versus x and the First order Phase Transition region (FPT) in Fig (8a),  (8b): The graph of $u(r)$ against x for free parameters and the FPT points,   (8c): comparing $r_{k}$ with respect to $x$ for($\alpha\rightarrow\gamma$) and ($\gamma\rightarrow\alpha$) and the coordinate of contact point   Fig (8d): $\Delta r_{k}$ with respect to $x$
@@
As evident from Fig(8d), just like before, at the initial of phase transition of the process is predominantly characterized by a direct transition. As observed in the diagram, the positive region is effectively the dominant area and no reactive structure appears until the end of the process.
@@
Figure 9: (9a): The $u(r)$ against x for free parameters and the FPT points,   (9b): comparing $r_{k}$ with respect to $x$ for($\alpha\rightarrow\gamma$) and ($\gamma\rightarrow\alpha$) and the coordinate of contact point   (9c):$\Delta r_{k}$ with respect to $x$
@@
As illustrated in Fig(9), with the progression of the phase transition process and as we approach the end, although the primary contribution is still from a direct transition, as depicted in Fig(9d), but the negative region has undergone significant growth and has become comparable to the positive region. It appears that, similar to the previous case, the reactive processes have become more potent in near the end of the process.
5 Discussion and comparison
In this section, for a better understanding, comparison, and interpretation of the results, it is first beneficial to delve a little into the concept of transition from small to large black holes (or large to small). The transition from a small to a large black hole can generally occur through the accretion of matter and energy from the surrounding environment of the black hole, or it can also occur through the merging of black holes, which may be accompanied by intensified emission of gravitational waves. However, from a thermodynamic perspective, this transition seems to encompass any physical process that facilitates changes in temperature, pressure, and other thermodynamic parameters in such a way that it leads to an increase in the radius of the event horizon and consequently, the enlargement of the black hole, which, given the super-gravitational structure of black holes, is a highly probable process. The reverse process, the transition from a large to a small black hole, can occur through processes such as Hawking radiation, where black holes emit particles due to quantum effects near the event horizon. Over time, this radiation can lead to a reduction in the mass and size of the black hole and its eventual evaporation. From a thermodynamic perspective, this phenomenon is equivalent to any process that creates a change in thermodynamic parameters leading to a reduction in the radius of the horizon and the shrinking of the black hole.
As observed in Figs (5b),(6a),(8b), (7a) and (9a), our black hole initially resides in a stable local minimum,$\alpha$, which globally is considered an unstable minimum. Consequently, as one would naturally expect, with the slightest thermal perturbation, and corresponding pressure, the system increasingly tends to transition towards the global minimum,$\gamma$, signifying an increase in radius.
@@ 5 Discussion and comparison
In both the NLM-C-Q-PFD and EGB-YM-CS models, upon examining the escape rate diagrams, we observe that once the radius exceeds a certain threshold, a small domain emerges where the reverse process forms and prevails. In other words, with the dominance of $r_{k}(\gamma\rightarrow\alpha)$ over $r_{k}(\alpha\rightarrow\gamma)$, a condition arises allowing black holes to transition from $\gamma$ to $\alpha$. Notably, over time and the sequence of frames, this seemingly negligible and very small region at the onset of phase transition gradually transforms into a significant area in the final frames. At first glance, the transition from $\gamma$ to $\alpha$ may seem less compatible with physical conditions because, within the same energy structure and without topological changes in the energy function (i.e., maintaining the gamma state as the global minimum and alpha as the local minimum), the system prefers to revert to the ( $\alpha$ ) state with a smaller radius, leading to a decrease in entropy and an increase in energy level. This situation can be termed ’improbable’ or ’less likely’ from a physical standpoint. However, on the other hand, from a Kramers’ perspective on the reverse process, it can be stated that particles situated in a potential well under thermal fluctuations can always escape with a specific probability coefficient (which is a function of the potential depth). Therefore, from the perspective of particles in the gamma potential, this movement is not considered an improbable phenomenon. Moreover, if we consider the transition from a smaller to a larger black hole as a general action, a natural reaction should form against it. The increase in radius and the consequent decrease in energy level is a recognized form of this reaction, leading to the reality that the escape rate, after reaching a peak, undergoes a decreasing trend until the probability becomes zero. But does this zero probability occur at any desired radius?

### v3

@@ III.0.1 Behavioral comparison of landscape free energy
Before any calculations, it is better first compare the various forms of free energy definitions, namely $G_{L}$, U, and L. Theoretically, from the canonical perspective the number of particles and mass should remain constant, consequently only the entropy-temperature and pressure-volume terms will play a role in the first law of thermodynamics. Additionally, given that in Landau’s free energy, $X$ can be considered equivalent to volume, it appears that ultimately, all three introduced forms for the free energy landscape must coincide. In Figs (2) and (2), for the model under study, we have plotted these three definitions once according to their general structure and once based on their dimensionless state. Practically, in terms of the shape structure, all three forms have the same trend, but there are also some visible differences, some of which may be attributed to computational precision. Moreover, since Landau’s free energy has dimensions of energy, we used the $G_{c}$ to render it dimensionless, which seems to have increased the degree of displacement, although the shape has been preserved. An intriguing point that comes to mind before concluding this discussion is that in statistical ensembles, such as the grand canonical ensemble where terms related to the chemical potential play a more prominent role, the complexities and structural differences of these three definitions may become more apparent compared to this scenario. Concerning the above discussion and as seen in Fig (2), all the shapes of the free energy landscape (with a little tolerance for displacement) will practically have the same shape structure in the canonical view. Accordingly, the choice and use of each do not seem to have an advantage over the others. For this reason, we will henceforth utilize them under the designation of the free energy landscape, without loss of generality.
Figure 2: Comparison of the Gibbs landscape free energy ($G_{L}$), thermal potential ($U$), and Landau free energy ($L$) in the diagram in two states, i.e., with dimension in Fig. (2) and dimensionless in Fig. (2)
Now, to be able to interpret the behavior of the energy function, we need to have the graph of temperature changes Fig. (3) and Gibbs free energy in terms of temperature Fig. (3).
@@
To better interpret the behavior of free energy, we initially turn to the swallowtail Fig. (3). In this figure, the line (A-N) represents the small black hole, and( C-J ) represents the large black hole, both of which are locally stable with positive heat capacity, and the line (N-C) with negative heat capacity represents the intermediate black hole, which is unstable. As a classical principle, we know that systems always tend to choose states with the least amount of energy. This principle also holds for the free energy landscape, meaning the system will move towards its minimum. In other words, the minima of the energy function correspond to stable states, and its maxima represent unstable states. Considering the overall shape in Fig. (4), it is also important to note that in a system with multiple local minima, each of these minima, although locally stable relative to their immediate vicinity, may act as unstable states globally relative to the primary minimum. Even if the system is initially in these states, over time it will progress towards the ultimate global minimum.
In Fig. (4), as is evident, the potential at point A has a global minimum, which belongs to the branch of the small black hole. Over time, with the increase in temperature, in Fig. (4), the local extrema B (minimum) and C (maximum) begin to form. However, they are still in the form of globally unstable states, and the system tends to maintain its position at the primary minimum on the branch of the small black hole. The situation in Fig. (4) is similar to (4), with the difference that the local extrema have fully formed. With a gradual increase in temperature, the system progresses towards the formation of a large and stable black hole, as shown in Fig. (4). Yet, the small black hole also exists in a stable state, indicating a situation where two equivalent global minima are created. This state corresponds to the isotherm line (H-I) in Fig. (3), which seemingly has a maximum on the branch of the intermediate black hole (I) and a global minimum (H). However, in reality, (H) is located on both the small and large black holes, meaning that a state is created where both the small and large black holes appear simultaneously and stably. In other words, a coexistence phase between the two stable black holes is formed. Further, the process of temperature increase will strengthen the large black hole and weaken the small black hole, Fig. (4) and (4). The point that can be mentioned at the end of this discussion is related to the second-order phase transition. When we study the system at a temperature higher than the critical temperature, Fig. (4), the system will follow a path similar to (I-H-J) in Fig. (3), and practically it is directly and continuously transferred from the small black hole to the large one.
@@
III.0.2 Kramer’s escape rate
@@
In Fig.(5), we approach a state at the onset of phase transition where the chaotic process of phase transition has formed due to primary reasons, meaning the $\gamma$ minimum has found its initial form (as shown in Fig.(5), although, it has not yet become a global minimum. The small black hole (in the local minimum potential $\alpha$) has begun its intense movement towards the large black hole, considering the formation of a secondary minimum and prevailing conditions, which is clearly observed in Fig.(5). The probability of transition $\alpha\rightarrow\gamma$ is significantly higher than the other transition, Fig.(5), which is why the dominant transfer seems to be entirely from $\alpha\rightarrow\gamma$. Since we are at the beginning of this transition, it seems that the big black hole has not yet formed, or in other words, it lacks enough particles for the formation of reverse processes. Therefore, as it is evident from the Fig.(5), the structure is under the complete control of the transition from a small black hole to a large one, and $\Delta r_{k}$ is completely positive.
@@
Figure 5: t versus x and the First order Phase Transition region (FPT) in Fig. (5),  (5): The graph of $u(r)$ against x for free parameters and the FPT points,   (5): comparing $r_{k}$ with respect to $x$ for($\alpha\rightarrow\gamma$) and ($\gamma\rightarrow\alpha$) and the coordinate of contact point   Fig. (5): $\Delta r_{k}$ ($r_{k}(\alpha\rightarrow\gamma)-r_{k}(\gamma\rightarrow\alpha)$) with respect to $x$
@@
Figure 6: (6): The $u(r)$ against x for free parameters and the FPT points,   (6): comparing $r_{k}$ with respect to $x$ for($\alpha\rightarrow\gamma$) and ($\gamma\rightarrow\alpha$) and the coordinate of contact point   (6):$\Delta r_{k}$ with respect to $x$   (6): The enlarged detail of the right end of the Fig (6)
In the subsequent stage, we examine a state midway through the phase transition, where the minimum $\gamma$ has assumed a relatively complete form (as depicted in Fig.(6), and although it has not yet become a global minimum, it is on par in depth with the alpha minimum. If we were to consider this statically and in a single-frame format, it would appear that the escape rate and the probability of movement from the local beta maximum in both directions (i.e., towards the small or the large black hole) should be equal. However, it is important to note that on one hand, the environmental conditions imposing the phase transition on the system, and on the other hand, the shrinking radius —which is somewhat contrary to normal physical laws— still, make the probability of transition $\alpha\rightarrow\gamma$ greater than the other transition, as shown in Fig.(6). For this reason, the dominant transfer seems to continue to be from $\alpha\rightarrow\gamma$. Nevertheless, as is evident in Fig. (6), a collision point has formed between the escape rate diagrams, and beyond this contact point, it is clear that the reverse process, albeit in a very small region at the end of the diagram, predominates over the direct process. It appears that as we approach the end of the phase transition, a negative region (Figs. (6) and (6) is forming which somehow, by reinforcing feedback processes, sets the stage for a change in the initial conditions and ultimately halts the phase transition.
@@
Ultimately, we turn our attention to a state near the end of the phase transition pathway, where the $\gamma$ has taken on a relatively complete shape (as illustrated in Fig.(7) and has become a global minimum. Although the environmental conditions that are imposing the phase transition on the system continue to make the probability of transition $\alpha\rightarrow\beta$ greater than the other transitions, as depicted in Fig.(7). However, as can be seen from Fig. (7), the coordinates of the contact point have moved significantly upwards, to the extent that the probability of the reverse process cannot be ignored in comparison to the direct process, and now a larger region is dominated by this process, as shown in Fig.(7). As we observed throughout this sequence, the closer we got to the end of the phase transition, the contact point advanced towards, increasing the probability of the reverse transition. It seems that this reverse process may, in some way, be the system’s reaction to prevent the phase transition process from becoming uncontrollable and the radial growth that might overshadow the stability of the black hole. It may also be hypothesized that the black hole, to prevent the phase transition process from becoming uncontrollable, instead of a direct and sudden transition from the small black hole to the large black hole (akin to a completely damped oscillation), performs this process in the form of a slowly damped, quasi-oscillatory manner.
@@
IV 4D AdS Einstein-gauss-bonnet-Yang-Mills black hole with a cloud of strings
@@
Since we tried to explain the contents as much as possible in the previous sections, considering the computational similarities, in this section, we will only mention the basic relationships, and we will bring the similar and unnecessary diagrams in Appendix A. The second selected model is the (EGB-YM-CS) black hole. So the metric of 4D AdS Einstein-gauss-bonnet-yang-mills black hole with a cloud of strings is as follows52 ,
@@
For the main quantities of this model i.e. mass $M$, Hawking temperature $T_{H}$, entropy $S$, pressure $P$ and volume $V$ we will have52 ,
@@
where q is the charge of the black hole, ”a” is related to the density of clouds of strings and comes from the energy-momentum tensor of the clouds of strings and $g$ 222 Gauss-Bonnet coupling constant is displayed with g instead of $\alpha$ to avoid mistakes with clouds of strings parameter is a positive Gauss-Bonnet coupling constant.
Figure 7: (7): The $u(r)$ against x for free parameters and the FPT points,   (7): comparing $r_{k}$ with respect to $x$ for($\alpha\rightarrow\gamma$) and ($\gamma\rightarrow\alpha$) and the coordinate of contact point   (7):$\Delta r_{k}$ with respect to $x$
IV.0.1 The landscape free energy
According to Eqs. (II.1), (II.1), (II.1) and (II.1) for Gibbs and Landau free energy function and thermal potential of this model, we will have,
@@ IV.0.1 The landscape free energy
In this model, we still use Eq.(9) to make the equations dimensionless.
IV.0.2 Kramer’s escape rate
@@
Figure 8: t versus x and the First order Phase Transition region (FPT) in Fig. (8),  (8): The graph of $u(r)$ against x for free parameters and the FPT points,   (8): comparing $r_{k}$ with respect to $x$ for($\alpha\rightarrow\gamma$) and ($\gamma\rightarrow\alpha$) and the coordinate of contact point   Fig. (8): $\Delta r_{k}$ with respect to $x$
@@
As evident from Fig.(8), just like before, at the initial of phase transition of the process is predominantly characterized by a direct transition. As observed in the diagram, the positive region is effectively the dominant area and no reactive structure appears until the end of the process.
@@
Figure 9: (9): The $u(r)$ against x for free parameters and the FPT points,   (9): comparing $r_{k}$ with respect to $x$ for($\alpha\rightarrow\gamma$) and ($\gamma\rightarrow\alpha$) and the coordinate of contact point   (9c):$\Delta r_{k}$ with respect to $x$
@@
As illustrated in Fig.(9), with the progression of the phase transition process and as we approach the end, although the primary contribution is still from a direct transition, as depicted in Fig.(LABEL:9d), but the negative region has undergone significant growth and has become comparable to the positive region. It appears that, similar to the previous case, the reactive processes have become more potent in near the end of the process.
V Discussion and comparison
@@
In this section, for a better understanding, comparison, and interpretation of the results, it is first beneficial to delve a little into the concept of transition from small to large black holes (or large to small). The transition from a small to a large black hole can generally occur through the accretion of matter and energy from the surrounding environment of the black hole, or it can also occur through the merging of black holes, which may be accompanied by intensified emission of gravitational waves. However, from a thermodynamic perspective, this transition seems to encompass any physical process that facilitates changes in temperature, pressure, and other thermodynamic parameters in such a way that it leads to an increase in the radius of the event horizon and consequently, the enlargement of the black hole, which, given the super-gravitational structure of black holes, is a highly probable process. The reverse process, the transition from a large to a small black hole, can occur through processes such as Hawking radiation, where black holes emit particles due to quantum effects near the event horizon. Over time, this radiation can lead to a reduction in the mass and size of the black hole and its eventual evaporation. From a thermodynamic perspective, this phenomenon is equivalent to any process that creates a change in thermodynamic parameters leading to a reduction in the radius of the horizon and the shrinking of the black hole. As observed in Figs. (5),(6),(8), (7) and (9), our black hole initially resides in a stable local minimum,$\alpha$, which globally is considered an unstable minimum. Consequently, as one would naturally expect, with the slightest thermal perturbation, and corresponding pressure, the system increasingly tends to transition towards the global minimum,$\gamma$, signifying an increase in radius. In both the NLM-C-Q-PFD and EGB-YM-CS models, upon examining the escape rate diagrams, we observe that once the radius exceeds a certain threshold, a small domain emerges where the reverse process forms and prevails. In other words, with the dominance of $r_{k}(\gamma\rightarrow\alpha)$ over $r_{k}(\alpha\rightarrow\gamma)$, a condition arises allowing black holes to transition from $\gamma$ to $\alpha$. Notably, over time and the sequence of frames, this seemingly negligible and very small region at the onset of phase transition gradually transforms into a significant area in the final frames. At first glance, the transition from $\gamma$ to $\alpha$ may seem less compatible with physical conditions because, within the same energy structure and without topological changes in the energy function (i.e., maintaining the gamma state as the global minimum and alpha as the local minimum), the system prefers to revert to the ( $\alpha$ ) state with a smaller radius, leading to a decrease in entropy and an increase in energy level. This situation can be termed ’improbable’ or ’less likely’ from a physical standpoint. However, on the other hand, from a Kramers’ perspective on the reverse process, it can be stated that particles situated in a potential well under thermal fluctuations can always escape with a specific probability coefficient (which is a function of the potential depth). Therefore, from the perspective of particles in the gamma potential, this movement is not considered an improbable phenomenon. Moreover, if we consider the transition from a smaller to a larger black hole as a general action, a natural reaction should form against it. The increase in radius and the consequent decrease in energy level is a recognized form of this reaction, leading to the reality that the escape rate, after reaching a peak, undergoes a decreasing trend until the probability becomes zero. But does this zero probability occur at any desired radius?